Complete combustion of a 0.30-mol sample of a hydrocarbon, CxHy, gives 1.20 mol of CO2 and 1.35 mol of H2O. The molecular formula of the original hydrocarbon is
Select one:
a. C4H9
b. C7H20
c. C3H5
d. C3H8
e. C6H6
1
Expert's answer
2020-04-10T13:35:51-0400

CxHy + (x+"{\frac {y} {4}}")O2 = xCO2 + "{\frac {y} {2}}"H2O

"{\frac {n(CO_2)} {n(C_xH_y)}}=x={\frac {1.20} {0.30}}=4"

x = 4

"{\frac {n(H_2O)} {n(C_xH_y)}}={\frac {y} {2}}={\frac {1.35} {0.30}}=4.5"

y = 9

Then the empirical formula of CxHy = C4H9, answer a

(Calculations are right, but i can't claim this answer as a correct one because hydrocarbon C4H9 doesn't exist)